Shift summary - Owl 08/30/99

Julie Whitmore (jaws@fnpxsr)
Mon, 30 Aug 1999 07:12:07 CDT

Shift Summary for OWL shift, 30/Aug/99

On Shift: John Shields, Julie

Hours of beam available at all intensities: 7.0
Hours of beam available at nominal intensity: 7.0
Hours of E832 data (special runs not included): 5.9
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time: DART fresh start;
Debug and swap Et board for CsI channel 1496.

========================================================================
Runs ended this shift:
Run 14374 - 6.48M events - BEAM_832 (HCC bit for 1496 dies; spill unknown)
Run 14375 - 3.22M events - BEAM_832 (HCC bit for 1496 dead)

Tapes brought to FCC this shift:
(including tapes for online split)

========================================================================
Accesses during shift: Swap Et board.

========================================================================
Overview of shift:
Steady beam. Data taking interrupted by channel 1496 Et problem.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14374 Inherit smooth E832 data taking.

~95 00:37 Pipeline errors. Try resetting pipeline.
kdisp shows that we've lost TE anode readout
from calorimeter (stream 1).
End run. Do DRT_FRESH_START.

14375 1 00:50 ktevana errors. KFREND.

01:00 Notice on kdisp that HCC bit for channel 1496 is
not illuminated. Check dynode signal at monitor.
It's solidly dead. Page Sydney to see if we can
run anyway. He defers to Ed. Page Ed who says
we should fix it.

01:30 Check HV is ok. It is. Set laser for HCC.
Channel 1496 is the only channel in which the
HCC bit doesn't respond. Sydney decides to come
in. He suggests we check signal at Et before
making blockhouse access. He suggests paging
JasonH as an extra pair of hands.

01:52 JS and JAWS go on access. Find healthy signal
going into Et board. Diagnose that it is a
blown input buffer. Luckily, Jason is on his way.

02:00 JasonH arrives. JCH and JAWS go on access to
swap Et board. Syndey arrives. He verifies
with kdisp that 1496 is fixed.

02:35 End run 14375.

14376 02:40 Start run 14376. Set laser for HCC. Everything
is fixed. Out of access. Beam arrives ~spill 5.

03:00 Syndey and Jason bicker like an old married couple
about online monitoring of HCC bit inefficiencies.
Search through previous run verifies that 1496
started to have trouble in run 14374.

04:00 JAWS finally starts to fill out run sheet and
encounters another anomoly. lsr#stability plot
wrt database (last 2 spills) is screwed up.
Plot wrt 1st spill is ok. After looking at the
plot and kdisp for ~15min, shift crew and Sydney
decide that the histogram probably is not being
filled properly.

04:50 Onsplit requests tape mount - QKA021.

135 05:35 Missing FFFF word, stream 5. KFREND.

05:40 Shift crew spots some problems with DC and muon.
Email Hogan and RickT.

07:00 Ed calls. We tell him about the PMT gain drift
laser reference problem. We look through the file
and find that the reference file is filled with 0s.
Luckily, he made a 2nd copy of the file that is
not corrupted. The file is copied over. The
PMT gain drift wrt DB plot should be fixed at next
run. Hand over run to Day shift crew.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

o if we lose beam for >= 40 minutes, and the DA works, do a laser scan.

o keep your eyes on channel 450 (see log book p. 103). If it remains stable,
and we do a laser calibration, maybe we don't need to swap the divider
(ask Ed). If it is unstable, page Julie and Ed.